### v1

# Non-Uniqueness of Metasurfaces for Wave Transformations

Abstract
We show that a large family of tensorial metasurfaces can be found that perform an identical wave transformation, showing that even when the conditions of reciprocity and passivity are imposed, there still remain many solutions to the design problem. As an example, we explore the case of a metasurface that rotates a single input polarization, showing we can parameterize the set of equivalent reciprocal metasurfaces in terms of a single complex parameter. Through allowing dissipation and gain within the response, the surface can have many different functionalities in the orthogonal polarization, opening up a new route for the design of multiplexed metasurfaces.
I Introduction
Metasurfaces are planar media containing an array of polarizable elements Chen et al. (2016), carefully chosen to perform a desired transformation of an incoming wave. Although this concept covers many areas of wave physics Assouar et al. (2018), here the focus is on metasurfaces for electromagnetic waves. Within such surfaces, the elements are typically spaced closer than the wavelength, allowing the characterization of the array in terms of an effective electromagnetic boundary condition Achouri and Caloz (2021). Whereas partially or totally transmitting metasurfaces are typically characterized in terms of sheet transition conditions Tretyakov (2003); Wu et al. (2018); Lebbe et al. (2023), or a polarizability Holloway et al. (2009); Zalus̆ki et al. (2016), impenetrable surfaces are usually characterized in terms of a surface impedance Senior and Volakis (1995); Tretyakov (2003); Maci et al. (2011). Illustrated in Fig. 1. This amounts to specifying the ratio of the electric and magnetic fields at all points on the surface. There are a variety of approaches to the design of such metasurfaces, controlling e.g. the local reflection phase over the surface Yu and Capasso (2014); Yu et al. (2011a); Chen et al. (2014a); Wang et al. (2019), or the coupling of surface waves to free–space radiation Fong et al. (2010). In this work, we consider impenetrable metasurfaces described in terms of a tensorial surface impedance and investigate the uniqueness of the surface parameters with respect to a desired transformation of the field.
Figure 1: Field transformations with tensorial metasurfaces: A metasurface characterized in terms of a tensorial surface impedance, $\mathrm{Z}_{s}$ can perform an arbitrary transformation of an incident electromagnetic wave. Here a normally incident wave with electric field aligned along the $x$–axis is transformed into a normally reflected wave with electric field at an angle $\theta$ to the $x$–axis, an example treated in Sec. IV. The right hand panels (a,b) illustrate the behaviour of a tensorial metasurface in terms of the induced current $\mathbf{J}_{\parallel}$ and the electric field eigenvectors, $\mathbf{E}_{\parallel,i}$, in terms of which (ci,cii) any incident field can be decomposed
Although the terminology of a ‘metasurface’ is little more than a decade old, the idea is a continuation of a large body of earlier work on frequency selective surfaces Munk (2005), which was revived when similar ideas began to be implemented in optics Yu et al. (2011b). Recent work has taken the metasurface idea significantly further forward. For example, through matching the electric and magnetic polarizability of the surface elements in a transmitting sheet, all reflection can be eliminated, resulting in so–called ‘Huygens’ metasurfaces Pfeiffer and Grbic (2013); Epstein and Eleftheriades (2016); Chen et al. (2018), which sculpt the field while maintaining its direction of propagation. More recently, more exotic metasurfaces have been considered, including non–Hermitian and PT–symmetric elements Fleury et al. (2014); Tapar and Kishen (2021); Y. et al. (2022), which contain carefully designed distributions of loss and/or gain, allowing for an unusual control over e.g. the polarization of the output field. Novel forms of surface excitations have also been discovered at junctions between metasurfaces Horsley and Hooper (2014); Bisharat and Sievenpiper (2017), and most recently non–reciprocal Lawrence et al. (2018) and space–time varying surface parameters have been considered Li et al. (2020); Wang et al. (2020), enabling both multiplexed functionality and spectral control.
As mentioned above, here we investigate the uniqueness of the tensorial surface impedance, $\mathrm{Z}_{s}$, when enforcing a single given transformation of the electromagnetic field. This is relevant to the problem of designing multiplexed—more specifically polarization multiplexed—metasurfaces, where two or more functionalities are imposed in the same design. Unless the surface properties are e.g. re–configurable Shirmanesh et al. (2020), it is well known that this is a difficult design problem, and there have been numerous previous approaches Capers et al. (2022); Dai et al. (2023); Murtaza Rind et al. (2024), some of which are summarized in Ref. Xu et al. (2021).
Here we show that, once a single field transformation has been specified at every point on the surface then there remains a great deal of freedom in the parameters determining the surface impedance. The greatest freedom occurs for non–reciprocal surfaces, but even reciprocal ones always contain a single free complex parameter. When we constrain the surface to be both passive and reciprocal, we find that through increasing dissipation within the surface, and thereby decreasing the amplitude of the transformed field, we, in a sense, increase the non–uniqueness of the metasurface. This shows that, through carefully increasing dissipation within the structure, we gain more freedom to find a metasurface that has a specific polarization multiplexed function.
II The Tensorial Impedance Boundary Condition
We can characterize the reflection of electromagnetic waves from any thin impenetrable metasurface in terms of an impedance boundary condition written in terms of a $2\times 2$ tensor valued surface impedance Senior and Volakis (1995); Tretyakov (2003), $\mathrm{Z}_{s}$. To understand the origin of this boundary condition, consider Ohm’s law within the metasurface relating the electric field parallel to a boundary, $\mathbf{E}_{\parallel}$, to the surface currents, $\mathbf{j}_{s}$, via the surface conductivity, $\sigma_{s}$. By relating the surface currents to the magnetic field via $\mathbf{j}_{s}=\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}$ where $\hat{\mathbf{n}}$ is the surface normal, we have an equation relating the surface electric and magnetic fields,
$\mathbf{j}_{s}=\sigma_{s}\cdot\mathbf{E}_{\parallel}=\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}.$ (1)
In general, the surface conductivity is a $2\times 2$ tensor, indicating anisotropy, i.e. the vectorial surface current may not be directly proportional to the surface electric field. For example, a surface composed of an array of thin parallel wires aligned with the $\hat{\mathbf{z}}$ direction would be characterized by the surface conductivity $\sigma_{s}=\sigma_{zz}\,\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}$, indicating that current can only flow along the axis of the wires and only flows in response to the $\hat{\mathbf{z}}$ component of the electric field. As a matter of convention, interfaces are characterized not in terms of the conductivity but in terms of the surface impedance, which is its inverse, $\mathrm{Z}_{s}=\sigma_{s}^{-1}$. This yields the impedance boundary condition,
$\mathbf{E}_{\parallel}=\mathrm{Z}_{s}\cdot\left(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}\right).$ (2)
Note that the component of the cycle averaged Poynting vector normal to the metasurface is given by $\langle S_{n}\rangle=\frac{1}{2}{\rm Re}[\hat{\mathbf{n}}\cdot(\mathbf{E}_{\parallel}\times\mathbf{H}_{\parallel}^{\star})]=-\frac{1}{2}{\rm Re}[\mathbf{E}_{\parallel}\cdot(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}^{\star})]$ and therefore the power locally entering the metasurface $\langle P_{n}\rangle=-\langle S_{n}\rangle$ is proportional to the Rayleigh quotient of the impedance matrix,
$\langle P_{n}\rangle=\frac{1}{2}{\rm Re}\left[\left(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}\right)^{\star}\cdot\mathrm{Z}_{s}\cdot\left(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}\right)\right].$ (3)
Given our assumption that the metasurface is impenetrable, a positive value for $\langle P_{n}\rangle$ indicates a dissipative response, whereas a negative value indicates amplification, or gain. Given that a lossless metasurface will have $\langle P_{n}\rangle=0$ for all possible incident fields, lossless surfaces have an impedance tensor that is ‘${\rm i}$’ times a Hermitian matrix (a so–called anti–Hermitian operator). In the general case, we can decompose the impedance into Hermitian and anti–Hermitian parts, $Z_{s}=H_{1}+{\rm i}\,H_{2}$, where $H_{1,2}$ are Hermitian matrices. The dissipated power is thus proportional to the Rayleigh quotient of $H_{1}$, implying that passive surfaces (i.e. those without gain) have a positive semi–definite form of $H_{1}$.
Another important constraint on the impedance tensor is provided by reciprocity. Representing the freedom to interchange sources and detectors, this implies the symmetry of the electromagnetic Green function Novotny and Hecht (2012) $G(\bm{x}_{1},\bm{x}_{2})=G^{\rm T}(\bm{x}_{1},\bm{x}_{2})$, which is equivalent to the symmetry of the operators in Maxwell’s equations. Reciprocity thus requires the surface impedance to be a symmetric tensor, $\mathrm{Z}_{s}=\mathrm{Z}_{s}^{\rm T}$.
III Field transformations with tensorial metasurfaces
An important application of these metasurfaces is in wave transformations, where a given incident wave is reflected into a specified radiation pattern, which in general can include both near and far–fields. Such transformations include the redirection, focusing, and absorption of an incident wave. One simple method for finding the impedance distribution that enacts a given transformation is to directly invert Eq. (2). To do this, we write our complex valued surface fields in terms of their magnitude, e.g. $|E|=(\mathbf{E}_{\parallel}\cdot\mathbf{E}_{\parallel}^{\star})^{1/2}$ and “direction” $\hat{\mathbf{e}}=\mathbf{E}_{\parallel}/|E|$,
$\mathbf{E}_{\parallel}$ $=|E|\hat{\mathbf{e}}$
$\mathbf{H}_{\parallel}$ $=|H|\hat{\mathbf{h}}$ (4)
Substituting this decomposition into the earlier expression for the impedance boundary condition (2) we obtain the requirement
$\mathrm{Z}_{s}\cdot(\hat{\mathbf{n}}\times\hat{\mathbf{h}})=\frac{|E|}{|H|}\hat{\mathbf{e}}$ (5)
which tells us that at each point on the surface, the impedance tensor must rotate and rescale the complex unit vector $\hat{\mathbf{n}}\times\hat{\mathbf{h}}$ so that it is parallel to the electric field unit vector, $\hat{\mathbf{e}}$. When our desired field transformation has $\hat{\mathbf{e}}$ everywhere parallel to $\hat{\mathbf{n}}\times\hat{\mathbf{h}}_{\parallel}$, this can be achieved with an isotropic metasurface, characterized in terms of a scalar impedance, $\mathrm{Z}_{s}=|E|/|H|$. This is possible in e.g. a $2D$ geometry, where one field component always lies out of the plane and the other always in the plane.
Nevertheless, in the general case a tensor valued impedance is required to satisfy (5). A simple solution is $\mathrm{Z}_{s}=(|E|/|H|)\,\hat{\mathbf{e}}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}$, where $\otimes$’ is the tensor product defined as e.g. $(\mathbf{a}\otimes\mathbf{b})_{ij}=\left(\mathbf{a}\mathbf{b}^{T}\right)_{ij}=a_{i}b_{j}$. However, as also explained in Ref. Kwon (2020), this form of surface impedance is most often neither passive ($\mathrm{Z}_{s}+\mathrm{Z}_{s}^{\dagger}$ not positive semi–definite), nor reciprocal ($\mathrm{Z}_{s}\neq\mathrm{Z}_{s}^{\rm T}$). This certainly represents a challenge for any practical realizations of these metasurface parameters, as both a power source and e.g. magnetic elements would have to be embedded within the surface! Yet there is an additional freedom that we can exploit to bring the design closer to reality. We can always perform a transformation $\mathrm{Z}_{s}\to\mathrm{Z}_{s}+\mathbf{v}\otimes\hat{\mathbf{h}}$, where $\mathbf{v}$ is an arbitrary vector, without affecting condition (5). This leaves us with an impedance tensor of the general form,
$\mathrm{Z}_{s}=\frac{|E|}{|H|}\left[\hat{\mathbf{e}}\otimes\left(\hat{\mathbf{n}}\times\hat{\mathbf{h}}\right)^{\star}+\mathbf{v}\otimes\hat{\mathbf{h}}\right].$ (6)
This surface impedance performs the desired transformation (5). Generally, it is non–reciprocal and non–passive, with the freedom to choose both complex components of $\mathbf{v}$, arbitrarily. The remainder of this paper explores this freedom in the design of metasurfaces.
III.1 Reciprocal metasurfaces:
To illustrate special cases of Eq. (6), we adopt a complex basis to express the impedance tensor. Here, we use the complex basis vectors $\hat{\mathbf{h}}$ and $\hat{\mathbf{n}}\times\hat{\mathbf{h}}^{\star}$ and the dual basis, $\hat{\mathbf{h}}^{\star}$ and $\hat{\mathbf{n}}\times\hat{\mathbf{h}}$, with the electric field direction being given by $\hat{\mathbf{e}}=e_{1}\,\hat{\mathbf{h}}+e_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}=\bar{e}_{1}\,\hat{\mathbf{h}}^{\star}+\bar{e}_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})$ and the arbitrary vector is written as $\mathbf{v}=\alpha\,\hat{\mathbf{h}}+\beta\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}=\bar{\alpha}\hat{\mathbf{h}}^{\star}+\bar{\beta}(\mathbf{n}\times\hat{\mathbf{h}})$. In terms of these basis vectors the impedance tensor (6) becomes
$\mathrm{Z}_{s}=\frac{|E|}{|H|}\bigg{\{}e_{1}\left[\,\hat{\mathbf{h}}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}+\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes\hat{\mathbf{h}}\right]+e_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}+\alpha\,\hat{\mathbf{h}}\otimes\hat{\mathbf{h}}\bigg{\}}$ (7)
where we have here ensured that the metasurface is reciprocal, $\mathrm{Z}_{s}=\mathrm{Z}_{s}^{\rm T}$, through taking
$\beta=e_{1},$ (8)
which determines a single component of the arbitrary vector, i.e. $\mathbf{v}\cdot(\hat{\mathbf{n}}\times\hat{\mathbf{h}})=\hat{\mathbf{e}}\cdot\hat{\mathbf{h}}^{\star}$.
From Eq. (7), we can see that, irrespective of the field transformation we seek to enact, it is always possible to find a reciprocal metasurface that will perform this transformation. Indeed, we have an infinite family of such surfaces as we are still free to choose the complex valued parameter $\alpha$. Provided we can implement amplification across the metasurface, this shows that there is a non-uniqueness in the surface parameters. The physical reason for this design freedom is that at every point we are imposing the response to a given surface field. Of course this leaves the response to the orthogonal surface field unspecified. The somewhat surprising fact is that while the requirement of reciprocity removes some of this redundancy, it is not enough to completely determine the form of the surface impedance.
III.2 Passive metasurfaces:
As described in Sec. II, for a metasurface to be passive it cannot amplify the incident field, and therefore the power entering the surface obeys $\langle P_{n}\rangle=-\langle S_{n}\rangle>0$. In terms of the unit vectors (III) and the basis introduced in the previous section, this is equivalent to the condition,
$\langle P_{n}\rangle={\rm Re}\left[\hat{\mathbf{e}}\cdot(\hat{\mathbf{n}}\times\hat{\mathbf{h}}^{\star})\right]={\rm Re}[\bar{e}_{2}]\geq 0.$ (9)
This is a requirement on the field transformation that we must implement at every point on the surface at the design stage: we cannot hope for a passive surface impedance if our design fields require the metasurface to generate energy! However, even assuming Eq. (9) holds, it may still be that the impedance (7) exhibits gain when probed with other incident fields (III).
As explained in Sec. II, to ensure the surface is passive, the Hermitian part of the impedance tensor must be positive semi–definite, i.e. for all complex unit vectors $\mathbf{u}$ we must have,
${\rm Re}\left[\mathbf{u}^{\star}\cdot\mathrm{Z}_{s}\cdot\mathbf{u}\right]$ $=\frac{|E|}{|H|}{\rm Re}\left[\bar{\alpha}|\bar{u}_{1}|^{2}+\bar{e}_{2}|\bar{u}_{2}|^{2}+\bar{e}_{1}\bar{u}_{1}^{\star}\bar{u}_{2}+\bar{\beta}\bar{u}_{2}^{\star}\bar{u}_{1}\right]$
$\geq 0.$ (10)
We can ensure the real part of the quadratic form (10) is always positive through imposing both condition (9) and choosing $\bar{\alpha}$, such that the determinant of the following Hermitian matrix is positive
${\rm det}\left[\left(\begin{matrix}{\rm Re}[\bar{\alpha}]&\frac{\bar{\beta}+\bar{e}_{1}^{\star}}{2}\\ \frac{\bar{\beta}^{\star}+\bar{e}_{1}}{2}&{\rm Re}[\bar{e}_{2}]\end{matrix}\right)\right]>0$ (11)
which implies the constraint
${\rm Re}[\bar{\alpha}]\geq\frac{1}{{\rm Re}[\bar{e}_{2}]}\left|\frac{\bar{e}_{1}^{\star}+\bar{\beta}}{2}\right|^{2}.$ (12)
Fig. 2 uses a randomly generated set of surface electric and magnetic fields to show that imposing this constraint indeed results in a passive metasurface. Note that if we impose zero dissipation in the response to the design field, ${\rm Re}[\bar{e}_{2}]=0$, the inequality (12) is impossible to fulfill. This shows that reducing the efficiency of the design (i.e. decreasing the reflected power) can help to find a set of metasurface parameters that are, overall passive. We shall return to this point again in the next section.
Imposing conditions (9) and (12) therefore ensures that our metasurface design is gain–free. If, on top of this, we also impose the constraint (8), we find a surface that is both passive and reciprocal. However, there is not enough information in Eqns. (8) and (12) to completely determine both complex parameters $\alpha$ and $\beta$. There thus remains a non–uniqueness in the design of even reciprocal, passive metasurfaces.
Figure 2: Randomly generated wave transformations: Using a set of $500$ randomly generated complex unit vectors, $\hat{\mathbf{e}}$ and $\hat{\mathbf{h}}$, representing different choices of wave transformation, we computed the reciprocal surface impedance given by Eq. (7) (with $|E|/|H|=1$, plus the constraint (9)), and the eigenvalues of $(\mathrm{Z}_{s}+\mathrm{Z}_{s}^{\dagger})/2$, which determine the dissipation/gain in the metasurface. (a) $\alpha=0$ for all designs, giving rise to eigenvalues of both signs, generally requiring a non–passive metasurface. (b) $\alpha$ constrained by Eq. (12) (here taking the equality), giving rise to eigenvalues of a positive sign only, showing that all surfaces can be made passive, as desired. (c) Corresponding values of $\alpha$ for the metasurfaces considered in panel (b), showing that in general this parameter must be complex valued.
III.3 A special class of reciprocal, passive metasurfaces:
A simple special case of the results given in Secs. III.1 and III.2 is where the unit vector $\hat{\mathbf{h}}$ defined in Eq. (III) is real valued. In this case, we no longer have to distinguish the vector basis from its dual, as defined in Sec. III.1, and barred and un–barred quantities become equal, i.e. $\bar{\alpha}=\alpha$, $\bar{\beta}=\beta$, $\bar{e}_{2}=e_{2}$, $\bar{e}_{1}=e_{1}$, etc. Therefore, the three conditions—(8), (9), and (12)–for a passive, reciprocal metasurface become simply
$\beta$ $=e_{1}$
${\rm Re}[e_{2}]$ $\geq 0$
${\rm Re}[\alpha]$ $\geq\frac{1}{{\rm Re}[e_{2}]}|e_{1}|^{2}$ (13)
which, as discussed above still leaves ${\rm Im}[\alpha]$ undetermined, indicating an infinite family of passive reciprocal metasurfaces that perform the same transformation.
As an initial example, consider a metasurface that completely absorbs incident waves when the electric field is aligned along the $x$ axis. For a surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$, the direction of the electric field is $\hat{\mathbf{e}}=\hat{\mathbf{x}}$, and the magnetic field is $\hat{\mathbf{h}}=-\hat{\mathbf{y}}$, with $|E|/|H|=\eta_{0}$ where $\eta_{0}$ is the free space impedance. In this case, $e_{1}=0$, $e_{2}=1$ and the reciprocal impedance (7) reduces to
$\mathrm{Z}_{s}=\eta_{0}\left(\hat{\mathbf{x}}\otimes\hat{\mathbf{x}}+\alpha\,\hat{\mathbf{y}}\otimes\hat{\mathbf{y}}\right).$ (14)
with all of the constraints in Eq. (13) fulfilled so long as ${\rm Re}[\alpha]>0$. Our design formulae therefore reproduces the expected result that a reciprocal, passive metasurface that perfectly absorbs $x$ oriented incident fields, must have the impedance component $\mathrm{Z}_{s,xx}$ equal to the free space impedance, and the orthogonal component, $\mathrm{Z}_{s,zz}$ must have a positive real part. Note that the imaginary part of $\alpha$ is unconstrained.
IV Example: Polarization Conversion
We now take a slightly more complex illustration of the family of reciprocal metasurfaces described by Eq. (7), although we retain the assumption of surface uniformity: we consider a polarization converting metasurface. A normally incident plane wave polarized along the $x$ axis impinges on a surface with surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$. The desired field is reflected with some real amplitude $\rho$, with a $\theta$ change in the polarization angle away from $\hat{\mathbf{x}}$, as illustrated in Fig. 1. The surface electric field equals,
Figure 3: Reflected power vs. polarization rotation angle $\theta$: The co– and cross–polarized reflectivities, $R_{ij}=|r_{ij}|^{2}$ (see appendix A) for different metasurfaces, as a function of the polarization rotation angle $\theta$ from Fig. 1. (a) $\alpha=0$ and $\rho=1$. As anticipated from Fig. 2, without constraining the value of $\alpha$, the designer polarization ($R_{xx}$ and $R_{yx}$) conserves power, but the secondary polarization ($R_{yy}$ and $R_{xy}$) is amplified. (b) As in panel (a), but for $\alpha=1\times 10^{12}$, showing that the surface approaches being passive for large real $\alpha$, as predicted by Eq. (18)
$\mathbf{E}$ $=\left[{\rm e}^{-\mathrm{i}k_{0}z}+r\cos{(\theta)}{\rm e}^{\mathrm{i}k_{0}z}\right]\hat{\mathbf{x}}+\rho\sin(\theta){\rm e}^{\mathrm{i}k_{0}z}\,\hat{\mathbf{y}}$
$\stackrel{{\scriptstyle z=0}}{{\to}}\hat{\mathbf{x}}+\rho\hat{\mathbf{m}},$ (15)
where we have introduced the unit vector, $\hat{\mathbf{m}}=\cos(\theta)\hat{\mathbf{x}}+\sin(\theta)\hat{\mathbf{y}}$. The previous sections show that the surface impedance enacting this transformation is characterized in terms of the electric and magnetic field unit vectors $\hat{\mathbf{e}}$ and $\hat{\mathbf{h}}$, which in this case are given by,
$\hat{\mathbf{e}}$ $=\frac{\hat{\mathbf{x}}+\rho\hat{\mathbf{m}}}{\sqrt{1+\rho^{2}+2\rho\cos(\theta)}}$
$\hat{\mathbf{h}}$ $=\frac{-\hat{\mathbf{y}}+\rho\hat{\mathbf{z}}\times\hat{\mathbf{m}}}{\sqrt{1+\rho^{2}-2\rho\cos(\theta)}}.$ (16)
In this example, the unit vectors (16) describing the field on the surface are purely real. Therefore, we can find a passive reciprocal metasurface using the results derived in Sec. III.3. The reciprocal surface impedance tensor is given by Eq. (7), with the vector components $e_{1}=\hat{\mathbf{e}}\cdot\hat{\mathbf{h}}$ and $e_{2}=\hat{\mathbf{e}}\cdot(\hat{\mathbf{n}}\times\hat{\mathbf{h}})$ equal to,
$e_{1}$ $=-\frac{2\rho\sin(\theta)}{\sqrt{(1+\rho^{2})^{2}-4\rho^{2}\cos^{2}(\theta)}}$
$e_{2}$ $=\frac{1-\rho^{2}}{\sqrt{(1+\rho^{2})^{2}-4\rho^{2}\cos^{2}(\theta)}}$ (17)
In terms of these two components, the constraints for a passive metasurface (13) are fulfilled if both $\rho^{2}\leq 1$ (i.e. the reflected wave does not contain a greater energy flux than the incident one) and,
${\rm Re}[\alpha]\geq\frac{1}{1-\rho^{2}}\frac{4\rho^{2}\sin^{2}(\theta)}{\sqrt{(1+\rho^{2})^{2}-4\rho^{2}\cos^{2}(\theta)}}$ (18)
with the imaginary part of $\alpha$ free.
Given the uniformity of the surface impedance in this case, we can analytically find the reflection from the surface defined by Eqns. (7), (16), and (IV) to verify the above predictions. The reflection amplitudes, $r_{xx}$, $r_{xy}$, $r_{yx}$, and $r_{yy}$ are defined such that the first index corresponds to the output polarization, and the second index to the input polarization. The expressions for these are given in appendix A. The power carried in each polarization channel is given by the square of these amplitudes, e.g. $R_{xx}=|r_{xx}|^{2}$ tells us the fraction of power reflected in the $x$ channel for incidence in the same channel. Fig. 3 shows both the results of a full wave simulation, and the reflection computed analytically, verifying both the reciprocity of the surface, $R_{xy}=R_{yx}$ and that incident $x$ polarized waves are reflected with the desired polarization. As shown in Fig. 3a, when $\alpha=0$, at designer incidence, the output power is conserved. Meanwhile for the secondary incident polarization, the surface reflects with increased power, even a divergent value when $\alpha=2$, shown in Fig. 4. For the surface to be passive (as in e.g. Chen et al. (2014b); Cheng et al. (2014); Feng et al. (2013); Huang et al. (2014); Yan et al. (2005); Yang et al. (2014); Zhu et al. (2013)), our constraint (18) shows that for $\rho=1$ (i.e. $100\%$ efficiency for $x$ polarized incidence), this is only possible when ${\rm Re}[\alpha]\to\infty$. Figs. 3b and 4 show this is indeed the case, with the amplification diminishing with increasing $\alpha$.
Figure 4: Polarization multiplexed functionality: for a $\theta=\pi/4$ polarization rotator (the design field incident with $x$ polarization), incident $y$ polarized waves can be transformed very differently, depending on the value of the complex parameter $\alpha$ in Eq. (7) for a reciprocal surface impedance. The above shows that, while the proportional of reflected power $P_{x}$ for $x$ polarized incidence is always unity, that for orthogonal incidence, $P_{y}$ is subject to either absorption or amplification, depending on the sign of ${\rm Re}[\alpha]$, with lasing–like behaviour at $\alpha=2$. Note that the maximum output power, $P_{\rm max}$, here found through an appropriate combination of $x$ and $y$ polarizations always exhibits gain except as $|\alpha|\to\infty$. This is consistent with Fig. 3 and constraint (18), as $\rho=1$, $\alpha$ must become infinite for the surface to be passive.
Figure 5: Real part of eigenvalues of $\mathrm{Z}_{s}$ when varying $\theta$ and $\alpha$: (a) Changing reflected polarization angle, $\theta$, at $\alpha=0$. As already indicated in Fig. 3, one eigenvalue always corresponds to loss and the other to gain within these metasurfaces. Note that as $\theta\rightarrow 90^{\circ}$, the eigenvalues tends towards ±$\pm\eta_{0}$ (i.e. a perfect absorber for the incident polarization and a infinite gain surface for the reflected polarization), denoted as black dashed lines. Note also that at $\theta=0$, both eigenvalues diverge. This does not indicate infinite gain, as such a large impedance mismatch with free space only leads to total reflection. (b) Eigenvalues for $100\%$ ($\rho=1$) and $50\%$ ($\rho=0.5$) efficient reflectors, as a function of $\alpha$. The dashed line shows the constraint (18) where the surface is passive, $\alpha>0.646$, above which the eigenvalues have a positive real part.
To demonstrate the validity of our inequality (18), we work in terms of the eigenvectors of the surface impedance tensor: those directions along which an applied electric field results in a parallel current density. As discussed in Sec. II, the power dissipated within the surface when the surface current $\hat{\mathbf{n}}\times\hat{\mathbf{H}}_{\parallel}$ is parallel to one of these eigenvectors $\mathbf{u}_{i}$ is, from Eq. (3),
$\langle P_{n,i}\rangle$ $=\frac{1}{2}{\rm Re}\left[\mathbf{u}_{i}^{\star}\cdot\mathrm{Z}_{s}\cdot\mathbf{u}_{i}\right]$
$=\frac{1}{2}\left|\mathbf{u}_{i}\right|^{2}{\rm Re}[z_{s,i}].$ (19)
where $\mathrm{Z}_{s}\mathbf{u}_{i}=z_{s,i}\mathbf{u}_{i}$. Therefore if ${\rm Re}[z_{s,i}]<0$ for either $i=1$ or $i=2$, then the surface cannot be passive. However this is only a necessary condition. In general, the dissipated power should be positive for a general superposition, $c_{1}\mathbf{u}_{1}+c_{2}\mathbf{u}_{2}$. This power equals:
$\langle P_{n}\rangle=|c_{1}|^{2}\langle P_{n,1}\rangle+|c_{2}|^{2}\langle P_{n,2}\rangle+\frac{1}{2}{\rm Re}\left[z_{s,1}\,c_{1}c_{2}^{\star}\mathbf{u}_{2}^{*}\cdot\mathbf{u}_{1}+z_{s,2}\,c_{1}^{\star}c_{2}\mathbf{u}_{1}^{*}\cdot\mathbf{u}_{2}\right]\geq 0.$ (20)
When the two eigenvectors are orthogonal, this reduces to $\langle P_{n}\rangle=|c_{1}|^{2}\langle P_{1}\rangle+|c_{2}|^{2}\langle P_{2}\rangle>0$, which leads to the requirement that the eigenvalues of $\mathrm{Z}_{s}$ should have a positive real part. In the current example, the surface impedance matrix is both real and symmetric (provided the free parameter $\alpha$ is real). Therefore its eigenvectors are orthogonal and we simply require both eigenvalues $z_{s,1}$, $z_{s,2}$ to have a positive real part. Fig. 5 shows the real part of the two eigenvalues of the impedance tensor (7) both as a function of the angle of polarization rotation and the parameter $\alpha$. In Fig. 5 we see that the surface indeed becomes passive once the inequality (18) is satisfied.
V Summary and Conclusions
Impenetrable metasurfaces are often characterized in terms of a scalar valued surface impedance, the value of which encodes both the phase and amplitude of the reflected field. However, as has been previously recognised, the use of a tensorial surface impedance—practically implemented via an array of anisotropic, polarizable elements—offers a much richer set of possible field transformations. Here we have explored the problem of designing tensorial metasurfaces for enacting arbitrary field transformations. We have investigated the uniqueness of the surface parameters for a given field transformation, when the metasurface is constrained to be passive, reciprocal, or both.
Our first finding is that it is possible to find a reciprocal metasurface, whatever the wave transformation we wish to perform. The tensorial surface impedance is given by Eq. (7) and there are an infinite number of such surfaces, parameterized in terms of a single complex number, $\alpha$. In general this family of metasurfaces will exhibit gain and loss for some incident fields, showing that through carefully implementing loss and/or gain within the surface, we may obtain multiplexed functionality. As an example, we showed in Sec. IV that a metasurface rotating $x$ polarized fields by an angle $\theta$ could simultaneously show a variety of different functionalities for $y$–polarized radiation, e.g. infinite amplification, absorption, or rotation.
Nevertheless, despite a large body of fascinating work (e.g. Fleury et al. (2014)) and recent experimental progress Tapar and Kishen (2021); Y. et al. (2022), metasurfaces with gain are much more difficult to implement than lossy ones. Our second finding is that, provided the complex parameter $\alpha$ satisfies the inequality (12), we can simultaneously ensure the metasurface is reciprocal and passive (without gain) for an arbitrary incident field. Surprisingly the surface parameters remain non–unique, and there remain an infinite family of surfaces that perform the same transformation. One important point is that the right hand side of the inequality (12) diverges when the initial design is made $100\%$ efficient (zero dissipation, all energy reflected). This actually reduces the design freedom, implying that the free parameter $\alpha$ must be taken to infinity in order that the surface be both lossless and reciprocal. Therefore, when the efficiency of the design is reduced (e.g. reducing the overall amplitude of the reflected field), we find that this increases the range of reciprocal, passive metasurfaces that can enact the transformation. This was demonstrated in the example of a polarization rotating metasurface, where Fig. 5b shows that reducing the efficiency immediately yields an increased range of passive reciprocal surfaces that would otherwise require ${\rm Re}[\alpha]\to\infty$. We have thus found that a careful choice of loss within the metasurface actually aids the design, allowing for polarization multiplexed functionality, at the price of reducing the efficiency.
Appendix A Reflection from a tensor impedance surface
Writing the components of the surface impedance tensor in terms of the Cartesian basis vectors, $\hat{\mathbf{x}}$and $\hat{\mathbf{y}}$ as,
$\mathrm{Z}_{s}=\left(\begin{matrix}Z_{xx}&Z_{xy}\\ Z_{yx}&Z_{yy}\end{matrix}\right)$ (21)
and assuming a wave at normal incidence with electric and magnetic fields given by ($A_{x}$ amplitude incident with $x$ polarization and $A_{y}$ amplitude incident with $y$ polarization),
$\mathbf{E}$ $=A_{x}\left[{\hat{\mathbf{x}}}\left({\rm e}^{-{\rm i}k_{0}z}+r_{xx}{\rm e}^{{\rm i}k_{0}z}\right)+{\hat{\mathbf{y}}}r_{yx}{\rm e}^{{\rm i}k_{0}z}\right]+A_{y}\left[{\hat{\mathbf{y}}}\left({\rm e}^{-{\rm i}k_{0}z}+r_{yy}{\rm e}^{{\rm i}k_{0}z}\right)+{\hat{\mathbf{x}}}r_{xy}{\rm e}^{{\rm i}k_{0}z}\right]$
$\eta_{0}\mathbf{H}$ $=A_{x}\left[{\hat{\mathbf{y}}}\left(-{\rm e}^{-{\rm i}k_{0}z}+r_{xx}{\rm e}^{{\rm i}k_{0}z}\right)-{\hat{\mathbf{x}}}r_{yx}{\rm e}^{{\rm i}k_{0}z}\right]+A_{y}\left[{\hat{\mathbf{x}}}\left({\rm e}^{-{\rm i}k_{0}z}-r_{yy}{\rm e}^{{\rm i}k_{0}z}\right)+{\hat{\mathbf{y}}}r_{xy}{\rm e}^{{\rm i}k_{0}z}\right],$ (22)
the impedance boundary condition defined above in Eq. (2) leads to the following expressions for the four reflections coefficients $r_{ij}$,
$r_{xx}$ $=\frac{(Z_{xx}-1)(Z_{yy}+1)-Z_{xy}Z_{yx}}{(Z_{xx}+1)(Z_{yy}+1)-Z_{xy}Z_{yx}}$ (23)
$r_{xy}$ $=\frac{2Z_{yx}}{(Z_{xx}+1)(Z_{yy}+1)-Z_{xy}Z_{yx}}$ (24)
$r_{yx}$ $=\frac{2Z_{xy}}{(Z_{xx}+1)(Z_{yy}+1)-Z_{xy}Z_{yx}}$ (25)
$r_{yy}$ $=\frac{(Z_{xx}-1)(Z_{yy}-1)-Z_{xy}Z_{yx}}{(Z_{xx}+1)(Z_{yy}+1)-Z_{xy}Z_{yx}}.$ (26)
These expressions we used to generate Figs. 35 in the main text.